OpenStudy (anonymous):

f(x)=ln(x)/7sqrt(x) Find the interval on which f is increasing. Find the interval on which f is decreasing. Find the local maximum value of f. Find the inflection point. Find the interval on which f is concave up. Find the interval on which f ins concave down. Please and thank you! @Mathematics

OpenStudy (anonymous):

That is; \[f(x)=\ln(x)/(7\sqrt{x})\] the square root x and 7 are BOTH under the ln(x)

myininaya (myininaya):

\[f(x)=\ln(\frac{x}{7 \sqrt{x}})\] like this?

OpenStudy (anonymous):

hints to you find f'(x) then f''(x)

myininaya (myininaya):

or do you mean what you actually typed which was \[f(x)=\frac{\ln(x)}{7 \sqrt{x}}\]

OpenStudy (anonymous):

increasing when f'(x) >0

OpenStudy (anonymous):

decreasing when f'(x)<0

OpenStudy (anonymous):

max when deriv = 0

OpenStudy (anonymous):

infl when f"(x) = 0 concave up when f"(x) >0 and down when f"(x)<0

OpenStudy (anonymous):

an interval is (#,#)... as in (-infinity,0), not just f(x)<0

OpenStudy (anonymous):

Everything you told me I know. The hard part is finding, for concave up for instance, where f''(x)>0...... so this doesn't help at all..

OpenStudy (anonymous):

anybody

OpenStudy (stacey):

Do you know how to find the derivative of the function you gave?